OpenStudy (anonymous):

Help please....4th grader needs help

OpenStudy (anonymous):

Yes?

OpenStudy (anonymous):

\[(3\frac{ 1 }{ 5}+1\frac{ 2 }{ 5})+4\frac{ 4 }{ 5 }\]

OpenStudy (anonymous):

9 2/5

OpenStudy (anonymous):

could you write out the steps how you got the answer please

OpenStudy (anonymous):

Add what's in parenthesis first. First whole numbers together, 3+1=4. Next add the two fractions together, remember to keep denominator same since your just adding, you get 3/5. Put them together and you get 4 3/5. Next add the 4 4/5. To do this add the whole numbers together, you get 8, and then add the numerators together, you'll get 7/5. Now since it's not proper to keep a fraction like this, subtract the denominator from the numerator,and add one to the whole number, whatever you get left over you'll put as your numerator. Thus, in the end you'll get 9 2/5 as you answer. Did this help at all?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

No problem, its just adding mixed numbers, and any time you need help just ask. :)

OpenStudy (anonymous):

thank you!! :)